# Characterization of Deletion/Substitution Channel Capacity for Small Deletion and Substitution Probabilities

### v1

@@ IV Random numbers of deletions and substitutions
Going from the case of fixed numbers of deletions and substitutions to the case where the number of deletions and substitutions are random merely involves taking advantage of the concentration of the numbers of deletions and substitutions around their means, $np_{d}$ and $n(1-p_{d})p_{s}$, respectively.
@@
In the setup with random numbers of deletions and substitutions, suppose there exists a decoding algorithm that succeeds with probability at least $\delta$ for codeword length of $n\geq\frac{12\log\frac{4}{\delta}}{\min\{p_{d},(1-p_{d})p_{s}\}}$. Then, the codebook size $N$ satisfies
@@
where $q_{d,\min}=(1-\gamma)np_{d}$, $q_{s,\min}=(1-\gamma)n(1-p_{d})p_{s}$, $q_{d,\max}=(1+\gamma)np_{d}$, $q_{s,\max}=(1+\gamma)n(1-p_{d})p_{s}$, with Let $\gamma=\sqrt{\frac{3\log\frac{4}{\delta}}{np_{d}}}$, and
@@
with
$t_{\max}^{*}=\lceil{3q_{d,\max}^{*}}\log\frac{ne}{q_{d,\max}^{*}}+{3q_{s,\max}^{*}}\log\frac{e(n-q_{d,\max}^{*})}{q_{s,\max}^{*}}+\log\frac{8}{\delta^{2}}\rceil.$ (26)
@@
Suppose we have a decoding algorithm for the deletion/substitution channel that succeeds on codebook $\mathcal{C}$ with probability $\delta\geq 0$. Standard multiplicative Chernoff gives an upper bound on the probability that a random variable lies outside an interval around its mean. More specifically (such as [18][Corollary 4.6]), if $X_{i}$’s are Bernoulli random variables with success probability $p_{i}$, then, $Pr(|\sum_{i}X_{i}-\mu|\geq\gamma\mu)\leq 2e^{-\frac{\mu\gamma^{2}}{3}}$, where $\mu=\sum_{i}p_{i}$. With $q_{d}=\sum_{i}X_{i}$, which is the random variable representing the number of deletions, we have $Pr(|q_{d}-np_{d}|\leq\gamma np_{d})\geq 1-2e^{-\frac{np_{d}\gamma^{2}}{3}}$. Setting the probability that the number of deletions, with mean $\mu_{d}=np_{d}$, lies in the interval $(1-\gamma)np_{d}\leq q_{d}\leq(1+\gamma)np_{d}$ to be at least $1-\frac{\delta}{2}$, we have $1-2e^{-\frac{np_{d}\gamma^{2}}{3}}=1-\frac{\delta}{2}$. This gives us $\gamma=\sqrt{\frac{3\log\frac{4}{\delta}}{np_{d}}}$. Finally, setting $\gamma\leq\frac{1}{2}$, as a design choice, we get, $n\geq\frac{12\log\frac{4}{\delta}}{np_{d}}$. Similarly, for the probability of the number of substitutions, with mean $\mu_{s}=n(1-p_{d})p_{s}$, lying in the interval $(1-\gamma)n(1-p_{d})p_{s}\leq q_{s}\leq(1+\gamma)n(1-p_{d})p_{s}$ to be at least $1-\frac{\delta}{2}$, we need $n\geq\frac{12\log\frac{4}{\delta}}{(1-p_{d})p_{s}}$. Combining the two conditions, we finally have $n\geq\frac{12\log\frac{4}{\delta}}{\min\{p_{d},(1-p_{d})p_{s}\}}$, as stated in the theorem statement.
@@
$\begin{split}&Pr\{(\text{Successful decoding})\\ &\hskip 28.45274pt\land(|q_{d}-\mu_{d}|\leq\gamma\mu_{d}\cap|q_{s}-\mu_{s}|\leq\gamma\mu_{s})\}\\ &\geq Pr\{\text{Successful decoding}\}\\ &\hskip 28.45274pt+Pr\{|q_{d}-\mu_{d}|\leq\gamma\mu_{d}\cap|q_{s}-\mu_{s}|\leq\gamma\mu_{s}\}-1\\ &=Pr\{\text{Successful decoding}\}\\ &\hskip 28.45274pt+Pr\{|q_{d}-\mu_{d}|\leq\gamma\mu_{d}\}Pr\{|q_{s}-\mu_{s}|\leq\gamma\mu_{s}\}-1\\ &\geq\frac{\delta^{2}}{4}.\end{split}$ (27)
Hence there must be some $(q_{d}^{*},q_{s}^{*})$ in the defined intervals such that the success probability of the exact $(q_{d}^{*},q_{s}^{*})$ deletion/substitution channel is at least $\frac{\delta^{2}}{4}$.
@@
Note that $\alpha^{*}$ is a non-decreasing function of both $q_{d}$ and $q_{s}$, hence is maximized with their maximum values. Finally, using the result in Theorem 1, by putting in each term the values in the intervals that maximize it, concludes the proof. ∎
@@
capacity of a deletion/substitution channel can be upper-bounded as
$C\leq 1-(1-o(1))(H(p_{d})+H(p_{s})).$ (30)
@@
Consider any fixed $p_{d}$ with $n$ going to infinity. We have $-n(1-\gamma)p_{d}\log((1-\gamma)p_{d})\leq\log{n\choose(1-\gamma)np_{d}}\leq-n(1-\gamma)p_{d}\log(\frac{1-\gamma}{e}p_{d})$. So $\log{n\choose(1-\gamma)np_{d}}=n(H(p_{d})+o(1))$, as $\gamma$ is vanishing. Similarly, for a any fixed $p_{s}$, as $p_{d}$ goes to zero, we have $\log{n-(1+\gamma)np_{d}\choose(1-\gamma)n(1-p_{d})p_{s}}=n(H(p_{s})+o(1))$.
Further, it can be shown that as the $\log\alpha_{\max}^{*}=\mathcal{O}(n(p_{d}+p_{d})\log(\log(1/p_{d})+\log(1/p_{s})))=o(n(H(p_{d})+H(p_{s}))$ as $p_{d}$ and $p_{s}$ go to 00. Therefore, the three terms $n-\log{n\choose(1-\gamma)np_{d}}-\log{n-(1+\gamma)np_{d}\choose(1-\gamma)n(1-p_{d})p_{s}}$ dominate the right-hand side of the equation; Dividing both sides by $n$, we obtain $\frac{1}{n}\log N\leq 1-(1-o(1))(H(p_{d})+H(p_{s}))$, which completes the proof. Note that the $o(1)$ term is understood as going to 0 in the limit as first $n$ goes to infinity and then as $p_{d}$ and $p_{s}$ go to 00. ∎
@@
Combined with the lower bound result in [18], it can be observed that the capacity of the deletion/substitution channel approaches $1-H(p_{d})-H(p_{s})$ as the probabilities of deletion $p_{d}$ and substitution $p_{s}$ go to zero.
@@ V Conclusion
In this paper, we characterized the capacity of binary channels with synchronization errors, where transmitted bits are randomly and independently deleted and flipped. To this end, we first obtained an upper bound on the capacity, assuming that the number of deletions and substitutions are known at the receiver. We then extended the results to the case of random numbers of deletions and substitutions. Finally, combining with an existing lower bound, we showed that, in the regime of asymptotically small deletion and substitution probabilities, the rate loss is equal to the sum of the binary entropies of the deletion and substitution probabilities; i.e., $C\rightarrow 1-H(p_{d})-H(p_{s})$.

### v2

@@ IV Random numbers of deletions and substitutions
Going from the case of fixed numbers of deletions and substitutions to the case where the numbers of deletions and substitutions are random merely involves taking advantage of the concentration of the numbers of deletions and substitutions around their means, $np_{d}$ and $n(1-p_{d})p_{s}^{\prime}$, respectively.
@@
In the setup with random numbers of deletions and substitutions, suppose that there exists a decoding algorithm that succeeds with probability at least $\delta$ for codeword length of $n\geq\frac{12\log\frac{4}{\delta}}{\min\{p_{d},(1-p_{d})p_{s}^{\prime}\}}$. Then, the codebook size $N$ satisfies
@@
where $q_{d,\min}=(1-\gamma)np_{d}$, $q_{s,\min}=(1-\gamma)n(1-p_{d})p_{s}^{\prime}$, $q_{d,\max}=(1+\gamma)np_{d}$, $q_{s,\max}=(1+\gamma)n(1-p_{d})p_{s}^{\prime}$, with Let $\gamma=\sqrt{\frac{3\log\frac{4}{\delta}}{np_{d}}}$, and
@@
with $t_{\max}^{*}=\lceil{3q_{d,\max}^{*}}\log\frac{ne}{q_{d,\max}^{*}}+{3q_{s,\max}^{*}}\log\frac{e(n-q_{d,\max}^{*})}{q_{s,\max}^{*}}+\log\frac{8}{\delta^{2}}\rceil$.
@@
Suppose that we have a decoding algorithm for the deletion/substitution channel that succeeds on codebook $\mathcal{C}$ with probability $\delta\geq 0$. The standard multiplicative Chernoff bound gives an upper bound on the probability that a random variable lies outside an interval around its mean. More specifically (such as [19][Corollary 4.6]), if $X_{i}$’s are Bernoulli random variables with success probability $p_{i}$, then, $Pr(|\sum_{i}X_{i}-\mu|\geq\gamma\mu)\leq 2e^{-\frac{\mu\gamma^{2}}{3}}$, where $\mu=\sum_{i}p_{i}$. With $q_{d}=\sum_{i}X_{i}$, which is the random variable representing the number of deletions, we have $Pr(|q_{d}-np_{d}|\leq\gamma np_{d})\geq 1-2e^{-\frac{np_{d}\gamma^{2}}{3}}$. Setting the probability that the number of deletions, with mean $\mu_{d}=np_{d}$, lies in the interval $(1-\gamma)np_{d}\leq q_{d}\leq(1+\gamma)np_{d}$ to be at least $1-\frac{\delta}{2}$, we have $1-2e^{-\frac{np_{d}\gamma^{2}}{3}}=1-\frac{\delta}{2}$. This gives us $\gamma=\sqrt{\frac{3\log\frac{4}{\delta}}{np_{d}}}$. Finally, setting $\gamma\leq\frac{1}{2}$, as a design choice, we get, $n\geq\frac{12\log\frac{4}{\delta}}{np_{d}}$. Similarly, for the probability of the number of substitutions, with mean $\mu_{s}=n(1-p_{d})p_{s}^{\prime}$, lying in the interval $(1-\gamma)n(1-p_{d})p_{s}^{\prime}\leq q_{s}\leq(1+\gamma)n(1-p_{d})p_{s}^{\prime}$ to be at least $1-\frac{\delta}{2}$, we need $n\geq\frac{12\log\frac{4}{\delta}}{(1-p_{d})p_{s}^{\prime}}$. Combining the two conditions, we finally have $n\geq\frac{12\log\frac{4}{\delta}}{\min\{p_{d},(1-p_{d})p_{s}^{\prime}\}}$, as stated in the theorem statement.
@@
$Pr\{(\text{Successful decoding})$
$\hskip 28.45274pt\land(|q_{d}-\mu_{d}|\leq\gamma\mu_{d}\cap|q_{s}-\mu_{s}|\leq\gamma\mu_{s})\}$
$\geq Pr\{\text{Successful decoding}\}$
$\hskip 28.45274pt+Pr\{|q_{d}-\mu_{d}|\leq\gamma\mu_{d}\cap|q_{s}-\mu_{s}|\leq\gamma\mu_{s}\}-1$
$=Pr\{\text{Successful decoding}\}$
$\hskip 28.45274pt+Pr\{|q_{d}-\mu_{d}|\leq\gamma\mu_{d}\}Pr\{|q_{s}-\mu_{s}|\leq\gamma\mu_{s}\}-1$
$\geq\frac{\delta^{2}}{4}.$ (24)
Hence, there must be some $(q_{d}^{*},q_{s}^{*})$ in the defined intervals such that the success probability of the exact $(q_{d}^{*},q_{s}^{*})$ deletion/substitution channel is at least $\frac{\delta^{2}}{4}$.
@@
Note that $\alpha^{*}$ is a non-decreasing function of both $q_{d}$ and $q_{s}$, hence it is maximized with their maximum values. Finally, using the result in Theorem 1, by putting in each term the values in the intervals that maximize it, concludes the proof. ∎
@@
The capacity of a deletion/substitution channel can be upper-bounded as
$C\leq 1-(1-o(1))(H(p_{d})+H(p_{s}^{\prime})).$ (27)
@@
Consider any fixed $p_{d}$ with $n$ going to infinity. We have $-n(1-\gamma)p_{d}\log((1-\gamma)p_{d})\leq\log{n\choose(1-\gamma)np_{d}}\leq-n(1-\gamma)p_{d}\log(\frac{1-\gamma}{e}p_{d})$. So $\log{n\choose(1-\gamma)np_{d}}=n(H(p_{d})+o(1))$, as $\gamma$ is vanishing. Similarly, for a any fixed $p_{s}$, as $p_{d}$ goes to zero, we have $p_{s}^{\prime}\rightarrow p_{s}$. So, we have $\log{n-(1+\gamma)np_{d}\choose(1-\gamma)n(1-p_{d})p_{s}^{\prime}}=n(H(p_{s})+o(1))$. Further, it can be shown that $\log\alpha_{\max}^{*}=\mathcal{O}(n(p_{d}+p_{d})\log(\log(1/p_{d})+\log(1/p_{s}^{\prime})))=o(n(H(p_{d})+H(p_{s}))$ as $p_{d}$ and $p_{d}$ (hence, also $p_{s}^{\prime}$) go to 00. Therefore, the three terms $n-\log{n\choose(1-\gamma)np_{d}}-\log{n-(1+\gamma)np_{d}\choose(1-\gamma)n(1-p_{d})p_{s}^{\prime}}$ dominate the right-hand side of the equation; Dividing both sides by $n$, we obtain $\frac{1}{n}\log N\leq 1-(1-o(1))(H(p_{d})+H(p_{s}))$, which completes the proof. Note that the $o(1)$ term is understood as going to 0 in the limit as first $n$ goes to infinity and then as $p_{d}$ and $p_{s}$ go to 00. ∎
@@
Combined with the lower bound result in [19], it can be observed that the capacity of the deletion/substitution channel approaches $1-H(p_{d})-H(p_{s})$ for the probabilities of deletion $p_{d}$ and substitution $p_{s}$ near zero.
@@ V Conclusion
In this paper, we characterized the capacity of binary channels with synchronization errors, where transmitted bits are randomly and independently deleted and flipped. To this end, we first obtained an upper bound on the capacity, assuming that the numbers of deletions and substitutions are known at the receiver. We then extended the results to the case of random numbers of deletions and substitutions. Finally, combining with an existing lower bound, we showed that, in the regime of asymptotically small deletion and substitution probabilities, the rate loss is equal to the sum of the binary entropies of the deletion and substitution probabilities. As a future work, it would be of interest to extend this result to channels with insertion, deletion, and substitution (IDS) errors.